OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. (Equation 1) x + y - 3z = -8(Equation 2) 2x + 2y + z = 12(Equation 3) 3x + y - z = -2. PLEASE HELP !

OpenStudy (anonymous):

now sovle the 1st and 2nd eqn and keeping in mind that z has to be removed so wwe have to make the coefficents of z equal: x+y-3z=-8 ---1eqn 6x+6y+3z=36---2 eqn after it was multipled by 3 so to make coefficents of z equal thenn adding these eqns we could remove z so the resulting eqn is: 7x+7y=28

OpenStudy (anonymous):

then solve like this for the 2nd and 3rd eqn by eliminating z....

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

glad to help :D !!!!!!!!